can you help me with this problem?
In 4 years, Nan will be 3 times as old as Dan will be then. The sum of their ages now is 56

Answer by oscargut(2103)

In 4 years, Nan will be 3 times as old as Dan will be then. The sum of their ages now is 56
hi,
Let x = age of Nan today
y = age of dan today

x+4 = 3(y+4)
x+y = 56
Then
x = 3y+12-4 = 3y+8
3y+8+y = 56
4y = 48
y = 12
x = 44
You can visit my web site:
http://www.online-maths-help.com/
Thanks